OpenStudy (anonymous):

Evaluate 4ab if a = -1 and b = 3. A. -12 B. -413 C. 12

OpenStudy (anonymous):

4ab=a*(-1)*(3)=-4*3=-12 Answer is C

OpenStudy (anonymous):

Whoops, I meant A